OpenStudy (anonymous):

Can someone please help pleaseeee!!! [7.02, 7.07] A greenhouse that specializes in growing vegetables used for Chinese cooking is divided into sections. The number of plants in each section depends on the number of sprinklers in that section. In a section with y sprinklers there are y(4y − 1) napa cabbage plants and (y + 5)2 bok choy plants. Which simplified expression represents the total number of napa cabbage plants and bok choy plants in a section with y sprinklers? 4y4 − 10y2 + 25 4y2 + 9y + 25 5y2 − 11y + 25 5y2 + 9y + 25

OpenStudy (anonymous):

wow they are really stretching to make these word problems. i think this means add \[ y(4y − 1)+ (y + 5)^2\]

OpenStudy (anonymous):

multiply first using the distributive property, then combine like terms

OpenStudy (anonymous):

oh thanks but like can u explain how to get the answer in detail?

OpenStudy (anonymous):

sure first of all \(y(4y-1)+(y+5)^2=y(4y-1)+(y+5)(y+5)\) then multiply \[y(4y-1)+(y+5)(y+5)=4y^2-y+y^2+5y+5y+25\] then combine like terms \[=5y^2+9y+25\]

OpenStudy (anonymous):

so yeah its just combining like terms thanks im going into proggraming so i need to know this

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

61